On January 2, 2015, Roth, Inc. purchased a laser cutting machine to be used in the fabrication of a part for one of its key products. The machine cost $130,000, and its estimated useful life was four years or 1,250,000 cuttings, after which it could be sold for $5,000. Required a. Calculate each year’s(2015-2019) depreciation expense for the machine's useful life under each of the following depreciation methods (round all answers to the nearest dollar): 1. Straight-line. 2. Double-declining balance. 3. Units-of-production. (Assume annual production in cuttings of 310,000; 460,000; 380,000; and 100,000.)
Answer #1

1.

Straight line Method
Year cost of equipment Less: Salvage value Depreciable value Depreciation
2015 $                   130,000 $                         5,000 $                   125,000 $125,000/4 = $31,250
2016 $                   130,000 $                         5,000 $                   125,000 $125,000/4 = $31,250
2017 $                   130,000 $                         5,000 $                   125,000 $125,000/4 = $31,250
2018 $                   130,000 $                         5,000 $                   125,000 $125,000/4 = $31,250

2.

Depreciation under DDB Method
Year - a Net Book value, beginning of year - b Double declained depreciation - c = b/Life of assets*2 Net book value, End of the year - d = b-c
2015 $                  130,000 $                                     65,000 $                     65,000
2016 $                    65,000 $                                     32,500 $                     32,500
2017 $                    32,500 $                                     16,250 $                     16,250
2018 $                    16,250 $16,250-$5,000 = $11,250 $                       5,000

3.

Year Units produced Depreciation (($130,000-$5,000) / 1,250,000 * units produced)
2015                 310,000 $                                                                                                                31,000
2016                 460,000 $                                                                                                                46,000
2017                 380,000 $                                                                                                                38,000
2018                 100,000 $                                                                                                                10,000

You can reach me over comment box if you have any doubts. Please rate this answer